# Quantum Inhomogeneous Field Theory: Unruh-Like Effects and Bubble Wall Friction

### v1

@@ Abstract
In this paper, we study a free scalar field in a specific (1+1)-dimensional curved spacetime. By introducing an algebraic state that is locally Hadamard, we derive the renormalized Wightman function and explicitly calculate the covariantly conserved quantum energy-momentum tensor up to a relevant order. From this result, we show that the Hadamard renormalization scheme, which has been effective in traditional quantum field theory in curved spacetime, is also applicable in the quantum inhomogeneous field theory. As applications of this framework, we show the existence of an Unruh-like effect for an observer slightly out of the right asymptotic region, as well as a quantum frictional effect on the bubble wall expansion during the electroweak phase transition in the early universe. Consequently, this study validates the consistency of our method for constructing meaningful physical quantities in quantum inhomogeneous field theory.
@@ 1 Introduction
where $m_{0}$ is a mass parameter, $\xi$ a dimensionless parameter, and $\mathcal{R}$ the curvature scalar of the background metric. The background curved spacetime $({\cal M},g_{ab})$ is not an arbitrary one but a specifically chosen by the supersymmetry requirement in our setup and therefore, the corresponding mass function is quite constrained. Then, the above scalar field action on ${\cal M}$ is converted to the scalar IFT action as
@@
which is our main interest in this paper. This action has been used for the bubble wall model in [13]. Specifically, we compute the pressure difference between slightly out of the Higgs phase region and the Higgs phase. At the end of the day, we show that the quantum effect of the vacuum for the observer who located near the Higgs phase lead to the frictional force against the bubble wall expansion.
This paper is organized as follows. In section 2, we review the supersymmetric formulation for our background and its corresponding IFT with canonical quantization. In this section, we also provide mode solutions and their corresponding various vacua. In section 3, after carefully introducing the ‘vacuum’ state for the observer residing slightly out of the asymptotic right region, we present the renormalized two-point functions for our model by adopting the so-called Hadamard renormalization method in the context of QFTCS. By employing the improved point-splitting method associated with the covariant energy-momentum tensor [34], we present our results on the VEV of energy-momentum tensor in our background in section 4. We provide concrete values of the VEV of energy-momentum tensor, consistent with the flat Minkowski limit. In section 5, we provide IFT interpretation of our results in section 5. Based on these results, we present the frictional quantum effects of our vacuum on the expansion of the bubble wall that occurs during the electroweak phase transition in the early universe. In the final section, we give some comments on our results and future directions.
@@ 2.1 SUSY background in (1+1) dimensions
along with the relation $\sqrt{-g}\,\mathcal{G}_{n}^{2}=-4\partial_{+}\mathcal{F}_{n}\partial_{-}\mathcal{F}_{n}$, where ±$\partial_{\pm}\equiv\frac{1}{2}\left(\frac{\partial}{\partial t}\pm\frac{\partial}{\partial x}\right)$. In the cases of flat and AdS2 spacetimes, the function $f$ becomes a constant, resulting in two supersymmetries corresponding to $\mathcal{N}=(1,1)$. However, when $f$ is not a constant and the Ricci scalar $\mathcal{R}$ depends only on the spatial coordinate $x$, there exists a single supersymmetry [25]. For further details, refer to [26].
@@
$\omega^{\prime\prime}+\frac{m_{0}}{2\xi}e^{\omega}\omega^{\prime}=0.$ (2.7)
@@ 2.2 Free SFTCS and supersymmetric IFT (SIFT)
It should be noted that since general covariance, which is a fundamental requirement in gravitational theory, is absent in the context of SIFT, all the physics in the SFTCS action (2.12) for different gauges can not be captured by the single SIFT action (2.16). What we have done is to convert the SFTCS action in a selected conformal gauge to the corresponding SIFT action.
@@
As seen from the form of the potential $V^{(1)}(x)=m_{\rm eff}^{2}(x)$ in (2.21), there are two sets of potentials determined by the sign of the solution parameter $a$, which is related to other parameters by $ab=\frac{m_{0}}{2\xi}$. For $a<0$, the geometry is divided by $x_{*}=-\frac{1}{b}\ln|a|$ into two asymmetric regions, $-\infty<x<x_{*}$ and $x_{*}<x<\infty$. In this scenario, the potential in SQM falls under the category of Eckart potentials. However, since we are interested in discussing the geometry defined over the entire spatial domain, $-\infty<x<\infty$, we focus on the case $a>0$. The two partner potentials $V^{(1)}$ and $V^{(2)}$ in SQM are then given by
$V^{(1)}$ $=2b^{2}\xi^{2}\left[1+\tanh\left(\frac{b}{2}(x-x_{0})\right)\right]-\frac{b^{2}\xi}{2}(2\xi-1)\frac{1}{\cosh^{2}\left(\frac{b}{2}(x-x_{0})\right)},$
$V^{(2)}$ $=2b^{2}\xi^{2}\left[1+\tanh\left(\frac{b}{2}(x-x_{0})\right)\right]-\frac{b^{2}\xi}{2}(2\xi+1)\frac{1}{\cosh^{2}\left(\frac{b}{2}(x-x_{0})\right)},$ (2.22)
which belong to the category of hyperbolic Rosen-Morse potentials. In fact, it can be shown that the two partner potentials in (2.2) satisfy the relation $V^{(2)}(x,\alpha_{1})=V^{(1)}(x,\alpha_{2})+R(\alpha_{1})$, where $\alpha_{2}$ is a function of $\alpha_{1}$. This means that the potential $V^{(1)}(x,\alpha_{1})$ transforms into the potential $V^{(2)}(x,\alpha_{1})$ by changing the parameter from $\alpha_{1}$ to $\alpha_{2}$ and adding a constant term $R(\alpha_{1})$. Using this shape invariance property of the two potentials, one can determine the spectrum and generalized eigenstates of the quantum mechanical system. We apply this method to find mode solutions of the scalar field $\phi$.
@@ 2.3 Mode solutions and canonical quantization
${\small V(x)=m^{2}_{\rm eff}(x)}$$x$$O$${\scriptstyle(b\beta)^{2}}$${\scriptstyle(b\beta-\epsilon)^{2}}$$x_{\epsilon}$
@@
$\phi_{\omega}(y)=(1+y)^{\gamma}\left[a_{1}y^{\alpha}F(A,B;C\,|\,-y)+a_{2}\,y^{\alpha+1-C}F(A-C+1,B-C+1;2-C\,|\,-y)\right],$ (2.23)
where $a_{1}$ and $a_{2}$ are integration constants, $y\equiv ae^{bx}=e^{b(x-x_{0})}$, $\alpha=\frac{i\omega}{b}$, $\gamma=\beta=2\xi=\frac{m_{0}}{ab}$, and
@@
We now quantize the scalar field $\phi$ using the mode solutions (2.23) and (2.3). First, we expand the scalar field using the mode solutions in (2.23) as follows:
@@
where $u^{(\mp)}_{\omega}(\boldsymbol{x})$ are given by
@@
The normalization is determined through the Klein-Gordon inner product. Note that
@@
Our normalization convention for $u^{(\mp)}_{\omega}(\boldsymbol{x})$ is consistent with the non-vanishing commutator of creation and annihilation operators, expressed as
$[a_{\omega}^{(i)},(a_{\omega^{\prime}}^{(j)})^{\dagger}]=\delta^{ij}\delta(\omega-\omega^{\prime}),$ (2.35)
where $i$ and $j$ denote -- or $+$. We refer to this canonical quantization as “L-quantization”. The operators $a^{(-)}_{\omega}$ and $a^{(+)}_{\omega}$ are not independent but proportional within the range $0\leq\omega\leq b\beta$, since no propagating degrees of freedom exist in the limit $x\to+\infty$. Thus, the commutation relation given in (2.35) is valid only for $\omega>b\beta$. Although the explicit commutation relations for $a_{\omega}^{(\mp)}$ in the range $0\leq\omega\leq b\beta$ are not provided, they are irrelevant for the following discussion. We then define the vacuum in L-quantization, $|0\rangle_{\rm L}$, as the state annihilated by $a_{\omega}^{(\mp)}$:
$a_{\omega}^{(\mp)}|0\rangle_{\rm L}=0.$ (2.36)
The Fock space constructed by $a^{(\mp)}_{\omega}$ and $\big{(}a^{(\mp)}_{\omega}\big{)}^{\dagger}$ is denoted by $\mathcal{F}_{\rm L}$.
In the left asymptotic region, $u^{(\mp)}_{\omega}(\boldsymbol{x})$ reduce to
@@
and in the L-quantization scheme, the form of the scalar field in the left asymptotic region $(x\to-\infty)$ approaches
$\phi_{\rm L}(\boldsymbol{x})\underset{x\to-\infty}{\simeq}\int^{\infty}_{0}\frac{d\omega}{\sqrt{2\pi}}\frac{1}{\sqrt{2\omega}}\left[a_{\omega}^{(+)}e^{-i\omega(t+x)}+a_{\omega}^{(-)}e^{-i\omega(t-x)}+\big{(}a_{\omega}^{(+)}\big{)}^{\dagger}e^{i\omega(t+x)}+\big{(}a_{\omega}^{(-)}\big{)}^{\dagger}e^{i\omega(t-x)}\right].$ (2.38)
The above equation indicates that L-quantization describes the system in terms of massless particles in the $x\to-\infty$ region. This description of massless particles is quite natural in the left asymptotic region but is not applicable in the right asymptotic region. Therefore, the L-quantization scheme is not suitable for the right asymptotic region or a right observer.
We now introduce another quantization scheme, referred to as “R-quantization.” From (2.3), we define $v^{(\mp)}_{k}(\boldsymbol{x})$ as
@@
$[b_{k}^{(i)},(b_{k^{\prime}}^{(j)})^{\dagger}]=\delta^{ij}\delta(k-k^{\prime}).$ (2.45)
@@
Now, we consider the field operator slightly out of the right asymptotic region. In the R-quantization scheme, the field operator can be expanded, based on the classical solution in (2.3), as
$\phi_{\rm R}(\boldsymbol{x})$ $=\int_{0}^{\infty}\frac{dk}{\sqrt{2\pi}}\frac{1}{\sqrt{2\omega}}\sum_{i=\pm}\left[b_{k}^{(i)}v^{(i)}_{k}(\boldsymbol{x})+\big{(}b_{k}^{(i)}\big{)}^{\dagger}\big{(}v^{(i)}_{k}(\boldsymbol{x})\big{)}^{*}\right]+\Phi_{\epsilon}(\boldsymbol{x}).$ (2.47)
Here, $\Phi_{\epsilon}(\boldsymbol{x})$ can be schematically written as
$\Phi_{\epsilon}(\boldsymbol{x})=\int_{b\beta-\epsilon}^{b\beta}\frac{d\omega}{\sqrt{2\pi}}\frac{1}{\sqrt{2\omega}}\left[c_{\omega}W_{\omega}(\boldsymbol{x})+c_{\omega}^{\dagger}(W_{\omega}(\boldsymbol{x}))^{*}\right],$ (2.48)
where we have neglected modes, which include exponentially decaying right-moving parts, in the energy range $0\leq\omega<b\beta-\epsilon$.
@@
Up to this point, we have discussed the canonical quantization of the scalar field from the perspective of IFT, which can also be understood as quantization within the FTCS framework for the right asymptotic region. This region corresponds to the weak curvature limit, where the FTCS framework is suitable for our analysis. Thus, R-quantization is valid within both the FTCS and IFT frameworks. However, L-quantization is not directly applicable to the FTCS framework due to the presence of a naked singularity as $x\to-\infty$.

### v4

@@ Abstract
In this paper, we study a free scalar field in a specific (1+1)-dimensional curved spacetime. By introducing an algebraic state that is locally Hadamard, we derive the renormalized Wightman function and explicitly calculate the covariantly conserved quantum energy-momentum tensor up to a relevant order. From this result, we show that the Hadamard renormalization scheme, which has been effective in traditional quantum field theory in curved spacetime, is also applicable in the quantum inhomogeneous field theory. As applications of this framework, we show the existence of an Unruh-like effect for an observer slightly out of the right asymptotic region, as well as the vanishing of quantum frictional effect in the leading order ($e^{-bx}$) on the bubble wall expansion during the electroweak phase transition in the early universe.
@@ 1 Introduction
[…]
@@
which is our main interest in this paper. This action has been used for the bubble wall model in [13]. Specifically, we compute the pressure difference between slightly out of the Higgs phase region and the Higgs phase. At the end of the day, we show that the quantum effect of the vacuum for the observer who located near the Higgs phase lead to the Unrhu-like effect and the vanishing of the frictional force in the leading order ($e^{-bx}$) against the bubble wall expansion.
This paper is organized as follows. In section 2, we review the supersymmetric formulation for our background and its corresponding IFT with canonical quantization. In this section, we also provide mode solutions and their corresponding various vacua. In section 3, after carefully introducing the ‘vacuum’ state for the observer residing slightly out of the asymptotic right region, we present the renormalized two-point functions for our model by adopting the so-called Hadamard renormalization method in the context of QFTCS. By employing the improved point-splitting method associated with the covariant energy-momentum tensor [34], we present our results on the VEV of energy-momentum tensor in our background in section 4. We provide concrete values of the VEV of energy-momentum tensor, consistent with the flat Minkowski limit. In section 5, we provide IFT interpretation of our results in section 5. Based on these results, we present the Unrhu-like effect and the vanishing of the frictional quantum effects of our vacuum in the leading order $(e^{-bx})$ on the expansion of the bubble wall that occurs during the electroweak phase transition in the early universe. In the final section, we give some comments on our results and future directions.
@@ 2.1 SUSY background in (1+1) dimensions
along with the relation $\sqrt{-g}\,\mathcal{G}_{n}^{2}=-4\partial_{+}\mathcal{F}_{n}\partial_{-}\mathcal{F}_{n}$, where $\partial_{\pm}\equiv\frac{1}{2}\left(\frac{\partial}{\partial t}\pm\frac{\partial}{\partial x}\right)$. In the cases of flat and AdS2 spacetimes, the function $f$ becomes a constant, resulting in two supersymmetries corresponding to $\mathcal{N}=(1,1)$. However, when $f$ is not a constant and the Ricci scalar $\mathcal{R}$ depends only on the spatial coordinate $x$, there exists a single supersymmetry [25]. For further details, refer to [26].
@@
$\Omega^{\prime\prime}+\frac{m_{0}}{2\xi}\Omega^{\prime}e^{\Omega}=0.$ (2.7)
@@ 2.2 Free SFTCS and supersymmetric IFT (SIFT)
It should be noted that since general covariance, which is a fundamental requirement in gravitational theory, is absent in the context of SIFT, all the physics in the SFTCS action (2.12) for different gauges can not be captured by the single SIFT action (2.16). What we have done is to convert the SFTCS action in a selected conformal gauge to the corresponding SIFT action. Therefore, the algebraic structures of quantum fields, including their causal structures, of FTCS and IFT (even in the absence of supersymmetry) are identical. Based on this observation, we proposed in [27] the application of the algebraic methodology developed in QFTCS to QIFT. This approach aligns with perspectives from researchers on algebraic quantum field theory, particularly in the context of CCR algebra (see, for example, Remark 8-(4) on page 14 of [4]).
@@
As seen from the form of the potential $V^{(1)}(x)=m_{\rm eff}^{2}(x)$ in (2.21), there are two sets of potentials determined by the sign of the solution parameter $a$, which is related to other parameters by $ab=\frac{m_{0}}{2\xi}$. For $a<0$, the geometry is divided by $x_{*}=-\frac{1}{b}\ln|a|$ into two asymmetric regions, $\infty<x<x_{*}$ and $x_{*}<x<\infty$. In this scenario, the potential in SQM falls under the category of Eckart potentials. However, since we are interested in discussing the geometry defined over the entire spatial domain, $\infty<x<\infty$, we focus on the case $a>0$. The two partner potentials $V^{(1)}$ and $V^{(2)}$ in SQM are given by
$V^{(1)}(x)$ $=2b^{2}\xi^{2}\left[1+\tanh\left(\frac{b}{2}(x-x_{0})\right)\right]-\frac{b^{2}\xi}{2}(2\xi-1)\frac{1}{\cosh^{2}\left(\frac{b}{2}(x-x_{0})\right)},$
$V^{(2)}(x)$ $=2b^{2}\xi^{2}\left[1+\tanh\left(\frac{b}{2}(x-x_{0})\right)\right]-\frac{b^{2}\xi}{2}(2\xi+1)\frac{1}{\cosh^{2}\left(\frac{b}{2}(x-x_{0})\right)}.$ (2.22)
These potentials belong to the category of hyperbolic Rosen-Morse potentials and satisfy the shape invariance condition. Here, $x_{0}$ is defined as $x_{0}\equiv-\frac{1}{b}\ln a$. It can be explicitly shown that the partner potentials in (2.2) satisfy the shape invariance relation
$V^{(2)}(x,\alpha_{1})=V^{(1)}(x,\alpha_{2})+R(\alpha_{2}),$ (2.23)
where $\alpha_{2}$ is a function of $\alpha_{1}$, i.e., $\alpha_{2}=\alpha_{2}(\alpha_{1})$. This implies that the potential $V^{(1)}(x,\alpha_{2})$ transforms into $V^{(2)}(x,\alpha_{1})$ by changing the parameter from $\alpha_{1}$ to $\alpha_{2}$ and adding a constant term $R(\alpha_{2})$.
More comments are given on the shape invariant property of the Rosen-Morse potentials in (2.2). General forms of the Rosen-Morse potential and its superpartner are given by [35]
$V^{(1)}(x;p,q,r)$ $=p^{2}+\frac{q^{2}}{p^{2}}+2q\tanh(rx)-\frac{p(p+r)}{\cosh^{2}(rx)},$
$V^{(2)}(x;p,q,r)$ $=p^{2}+\frac{q^{2}}{p^{2}}+2q\tanh(rx)-\frac{p(p-r)}{\cosh^{2}(rx)},$ (2.24)
where $p$, $q$, and $r$ are constant parameters. The shape invariant property in (2.23) for the general forms in (2.2) is given by
$V^{(2)}(x;p,q,r)=V^{(1)}(x;p-r,q,r)+R(p-r,q,r),$ (2.25)
where
$R(p-r,q,r)=p^{2}-\big(p-r\big)^{2}+q^{2}\Bigg(\frac{1}{p^{2}}-\frac{1}{(p-r)^{2}}\Bigg).$ (2.26)
The potentials in (2.2) are obtained from those in (2.2) by setting
$p=-b\xi,\quad q=b^{2}\xi^{2},\quad r=\frac{b}{2}.$ (2.27)
Then, the relation in (2.25) is also satisfied with this setting.
Using this shape invariance property of the two potentials, one can determine the spectrum and generalized eigenstates of the quantum mechanical system. We apply this method to find mode solutions of the scalar field $\phi$.
@@ 2.3 Mode solutions and canonical quantization
[…]
@@
$\phi_{\omega}(y)=(1+y)^{\beta}\left[a_{1}y^{\alpha}F(A,B;C\,|\,-y)+a_{2}\,y^{\alpha+1-C}F(A-C+1,B-C+1;2-C\,|\,-y)\right],$ (2.28)
where $a_{1}$ and $a_{2}$ are integration constants, $y\equiv ae^{bx}=e^{b(x-x_{0})}$, $\alpha=\frac{i\omega}{b}$, $\beta=2\xi=\frac{m_{0}}{ab}$, and
@@
We now quantize the scalar field $\phi$ using the mode solutions (2.28) and (2.3). First, we expand the scalar field using the mode solutions in (2.28) schematically as follows:
@@
where $u^{(\mp)}_{\omega}(\boldsymbol{x})$ are given by [^2] We define $a\equiv e^{-bx_{0}}$ and perform a change of coordinates $x-x_{0}\mapsto x$. In this new coordinate system, the Ricci scalar in (2.10) is given by $\mathcal{R}=2a^{2}b^{2}e^{-bx}.$
@@
Note that
@@
In principle, the normalization is determined through the Klein-Gordon inner product. If the mode functions (2.36) and (2.37) were correctly orthonormalized, our normalization convention for $u^{(\mp)}_{\omega}(\boldsymbol{x})$ would be consistent with the non-vanishing commutator of the creation and annihilation operators, expressed as
$a_{\omega}^{(i)},(a_{\omega^{\prime}}^{(j)})^{\dagger}]=\delta^{ij}\delta(\omega-\omega^{\prime}),$ (2.40)
where $i$ and $j$ denote - or ++.
However, the operators $a^{(-)}_{\omega}$ and $a^{(+)}_{\omega}$ are not independent but proportional in the range $0\leq\omega<b\beta$, due to the absence of propagating degrees of freedom in the limit $x\to+\infty$. Consequently, the commutation relation given in (2.40) is valid only for $\omega\gg b\beta$. Mode expansion and quantization become involved in the low-frequency regime, resulting from reflecting modes from the mass function wall within the range $0\leq\omega<b\beta$. To address this issue, a proposal has been made in the context of the spectral method [13], in which the mode functions can be properly normalized without relying on the Klein-Gordon inner product. According to the spectral method, the field expansion (2.35) should be rewritten in terms of new operators and new mode functions constructed by [^3] In fact, this construction is indirect and uses a mathematical tool known as a Weyl-Titchmarsh-Kodaira theory [13]. linear combinations of $u^{(i)}_{\omega}(\boldsymbol{x})$ and $(u^{(i)}_{\omega}(\boldsymbol{x}))^{*}$. Before going ahead, let us elaborate on the asymptotic behavior of the schematic expression in (2.35). In the left asymptotic region $(x\to-\infty)$, the mode functions $u^{(\mp)}_{\omega}(\boldsymbol{x})$ reduce to
@@
and the field operator in equation (2.35) approaches
$\phi_{\rm L}(\boldsymbol{x})$ $\underset{x\to-\infty}{\simeq}\int^{\infty}_{b\beta}\frac{d\omega}{\sqrt{2\pi}}\frac{1}{\sqrt{2\omega}}\Big[a_{\omega}^{(+)}e^{-i\omega(t+x)}+a_{\omega}^{(-)}e^{-i\omega(t-x)}+h.c.~\Big]$
$\qquad\quad+\int^{b\beta}_{0}\frac{d\omega}{\sqrt{2\pi}}\frac{1}{\sqrt{2\omega}}\Big[c_{\omega}\,e^{-i\omega(t+x)}+r_{\omega}c_{\omega}\,e^{-i\omega(t-x)}+h.c.~\Big]\,$, (2.42)
where $c_{\omega}$ is a new operator alluded above, $r_{\omega}$ denotes a certain phase factor, and $h.c.$ means the Hermitian conjugate. This expression reveals that the quantization describes the system in terms of massless particles in the $x\to$-$\infty$ region. As highlighted earlier, this field quantization deviates from conventional canonical quantization due to the interdependence of $u^{(-)}_{\omega}$ and $u^{(+)}_{\omega}$ in the low-frequency regime. We have coined this quantization scheme “L-quantization” in [26]. Within the L-quantization framework, the vacuum state $|0\rangle_{\rm L}$ is defined as the state annihilated by $a_{\omega}^{(\mp)}$ and $c_{\omega}$:
$a_{\omega}^{(\mp)}|0\rangle_{\rm L}=c_{\omega}|0\rangle_{\rm L}=0\,.$ (2.43)
In this study, our analysis focuses solely on the right asymptotic region. We defer a detailed investigation of L-quantization in the left asymptotic region to future research.
While the above L-quantization would be appropriate for the left asymptotic region, it does not extend to the right asymptotic region, rendering the quantization scheme unsuitable for that domain or for an observer situated there. We now introduce another quantization scheme, referred to as “R-quantization.” From (2.3), we define $v^{(\mp)}_{k}(\boldsymbol{x})$ as
@@
$b_{k}^{(i)},(b_{k^{\prime}}^{(j)})^{\dagger}]=\delta^{ij}\delta(k-k^{\prime}).$ (2.50)
@@
Now, we consider the field operator slightly out of the right asymptotic region. Proper normalization of the mode functions defined in (2.44) and (2.45) via the Klein-Gordon inner product is generally difficult to achieve, as L-quantization in the left region. As an alternative, the spectral method outlined earlier can be employed to obtain a properly normalized quantization, yielding results consistent with the L-quantization scheme. Nevertheless, as demonstrated below, within the range of approximations adopted in this study, the R-quantization based solely on the mode functions $v^{(i)}_{k}$ and $(v^{(i)}_{k})^{\dagger}$ provides consistent and reliable outcomes. Specifically, for the R-quantization slightly out of the right asymptotic region, the field operator can be expressed in terms of the mode functions introduced in (2.44) and (2.45) as follows:
$\phi_{\rm R}(\boldsymbol{x})$ $=\int_{0}^{\infty}\frac{dk}{\sqrt{2\pi}}\frac{1}{\sqrt{2\omega}}\sum_{i=\pm}\left[b_{k}^{(i)}v^{(i)}_{k}(\boldsymbol{x})+\big(b_{k}^{(i)}\big)^{\dagger}\big(v^{(i)}_{k}(\boldsymbol{x})\big)^{*}\right]+\Phi_{\epsilon}(\boldsymbol{x}),$ (2.52)
where $\Phi_{\epsilon}(\boldsymbol{x})$ accounts for the exponentially decaying modes as $x\to\infty$ in the energy range $0\leq\omega<b\beta$. In this paper, the main focus is physics at the position $x=x_{\epsilon}$ slightly out of the right asymptotic region (See Fig.1). One may worry about the contribution of propagating modes near the position $x=x_{\epsilon}$ over the range $b\beta-\epsilon\leq\omega<b\beta$. To give an order estimate near $x=x_{\epsilon}$, let us write (2.3) with $W(\boldsymbol{x})\sim e^{-i(\omega t\pm kx)}$ schematically as
$\Phi_{\epsilon}(\boldsymbol{x})=\int_{b\beta-\epsilon}^{b\beta}\frac{d\omega}{\sqrt{2\pi}}\frac{1}{\sqrt{2\omega}}\left[c_{\omega}W_{\omega}(\boldsymbol{x})+c_{\omega}^{\dagger}(W_{\omega}(\boldsymbol{x}))^{*}\right]+{\cal O}(e^{-bx_{\epsilon}})\,.$ (2.53)
where the exponentially decaying part comes from the range $0\leq\omega<b\beta-\epsilon$. As discussed in Appendix B, the first integral term over the range $b\beta-\epsilon<\omega<b\beta$ contributes at a higher order. Then, we can safely neglect the contribution of $\Phi_{\epsilon}$, which is ${\cal O}(e^{-2bx_{\epsilon}})$, to 2-point function $\langle\phi\phi\rangle$, since our analysis of the 2-point function $\langle\phi\phi\rangle$ focuses on terms up to ${\cal O}(e^{-bx_{\epsilon}})$.
@@
Up to this point, we have discussed the canonical quantization of the scalar field from the perspective of IFT, which can also be understood as quantization within the FTCS framework for the right asymptotic region. This region corresponds to the weak curvature limit, where the FTCS framework is suitable for our analysis. Thus, R-quantization is valid within both the FTCS and IFT frameworks. However, L-quantization is not directly applicable to the FTCS framework due to the presence of a naked singularity as $x\to$-$\infty$. We leave the issue of the quantization and its implications near the right asymptotic region for future investigation [39].